# Discovering Large-Scale Structure at in the C3VO Survey

### v1

@@ Acknowledgments
Some of the material presented in this paper is supported by the National Science Foundation under Grant Nos. 1411943 and 1908422. This work was partially supported by NASA’s Astrophysics Data Analysis Program under grant number 80NSSC21K0986. This work was additionally supported by the France-Berkeley Fund, a joint venture between UC Berkeley, UC Davis, and le Centre National de la Recherche Scientifique de France promoting lasting institutional and intellectual cooperation between France and the United States. This study is based, in part, on data products from observations made with ESO Telescopes at the La Silla Paranal Observatory under ESO programme ID 179.A-2005 and on data products produced by TERAPIX and the Cambridge Astronomy Survey Unit on behalf of the UltraVISTA consortium. This study is based, in part, on data collected at the Subaru Telescope and obtained from the SMOKA, which is operated by the Astronomy Data Center, National Astronomical Observatory of Japan. This work is based, in part, on observations made with the Spitzer Space Telescope, which is operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. UKIRT is supported by NASA and operated under an agreement among the University of Hawaii, the University of Arizona, and Lockheed Martin Advanced Technology Center; operations are enabled through the cooperation of the East Asian Observatory. When the data reported here were acquired, UKIRT was operated by the Joint Astronomy Centre on behalf of the Science and Technology Facilities Council of the U.K. This study is also based, in part, on observations obtained with WIRCam, a joint project of CFHT, Taiwan, Korea, Canada, France, and the Canada-France-Hawaii Telescope which is operated by the National Research Council (NRC) of Canada, the Institut National des Sciences de l’Univers of the Centre National de la Recherche Scientifique of France, and the University of Hawai’i. Supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. Some portion of the spectrographic data presented herein was based on observations obtained with the European Southern Observatory Very Large Telescope, Paranal, Chile, under Large Programs 070.A-9007 and 177.A-0837. The remainder of the spectrographic data presented herein were obtained at the W.M. Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W.M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
@@ Appendix A Boundary Effects in the Lightcone
Because of the finite size of the cubes that are used to construct the lightcones (see §4.1 for details), there exist boundary effects at the interface between cubes that have the potential to cause underrichness in structures whose centers lie near the edge of the cubes. In addition, the lightcone is constructed in such a way that it subtends a circular region of radius of 2.3 deg, which causes additional boundary effects at the border of the lightcone along the transverse dimensions. In order to determine what structures are affected by these boundary effects, we utilize a version of the lightcone that probes 2 deeper than the version of lightcone that is used for the vast majority of our mocks (i.e., IRAC125.5 vs. IRAC124.8) in order to search for simulated protostructures that are significantly underrich in member galaxies relative to the population average. A deeper IRAC1 cut is used for this exercise in order to differentiate protostructures that are underrich due to border effects from those that appear underrich in the IRAC124.8 due to a shift in the luminosity function of member galaxies to fainter luminosities for astrophysical reasons.
As discussed in §4.3, we break the sample of simulated protostructures into redshift bins of over the redshift range and perform a (log) linear fit of vs. with iterative clipping. All protostructures which are offset from the fit relation by 3 on the low end of this mass richness relation are removed from our sample and are not considered in any of the metrics presented in this paper. More specifically, we do not use any subsumed structures that contain a clipped protostructure in the mass fit. Formally, we should also remove these structures from consideration in the estimate of the completeness and the purity calculation as well, but the inclusion of these structures have a negligible effect on the overall completeness numbers and would introduce a great deal of complexity into the purity calculation for likely the same result.
In total, the clipped protostructures comprise only 0.6% of the total protostructure catalog. Note that this process will also clip protostructures that are deficient in galaxies brighter than IRAC125.5, which will potentially induce some bias in our results as, e.g., our completeness numbers may be overestimated for massive protostructures that are observed to be underrich in galaxies (e.g., Newman et al., 2022). However, such protostructures appear, at least in our lightcones to be rare (see also the simulation results presented in Newman et al. 2022). Figure 16 shows the -richness relation for all simulated protostructures as well as those that are identified as artificially underrich. Also plotted are the RA, Dec, and redshift histogram of all simulated protostructures along with those that are identified as artificially underrich.

### v2

@@ Acknowledgments
We thank the anonymous referee for their valuable comments in improving this work. Some of the material presented in this paper is supported by the National Science Foundation under grant Nos. 1411943 and 1908422. This work was partially supported by NASA’s Astrophysics Data Analysis Program under grant No. 80NSSC21K0986. This work was additionally supported by the France-Berkeley Fund, a joint venture between UC Berkeley, UC Davis, and le Centre National de la Recherche Scientifique de France promoting lasting institutional and intellectual cooperation between France and the United States. This study is based, in part, on data products from observations made with ESO Telescopes at the La Silla Paranal Observatory under ESO program ID 179.A-2005 and on data products produced by TERAPIX and the Cambridge Astronomy Survey Unit on behalf of the UltraVISTA consortium. This study is based, in part, on data collected at the Subaru Telescope and obtained from the SMOKA, which is operated by the Astronomy Data Center, National Astronomical Observatory of Japan. This work is based, in part, on observations made with the Spitzer Space Telescope, which is operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. UKIRT is supported by NASA and operated under an agreement among the University of Hawaii, the University of Arizona, and Lockheed Martin Advanced Technology Center; operations are enabled through the cooperation of the East Asian Observatory. When the data reported here were acquired, UKIRT was operated by the Joint Astronomy Centre on behalf of the Science and Technology Facilities Council of the U.K. This study is also based, in part, on observations obtained with WIRCam, a joint project of CFHT, Taiwan, Korea, Canada, France, and the Canada–France–Hawaii Telescope which is operated by the National Research Council (NRC) of Canada, the Institut National des Sciences de l’Univers of the Centre National de la Recherche Scientifique of France, and the University of Hawai’i. This paper was supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. Some portion of the spectrographic data presented herein was based on observations obtained with the European Southern Observatory Very Large Telescope, Paranal, Chile, under Large Programs 070.A-9007 and 177.A-0837. The remainder of the spectrographic data presented herein were obtained at the W.M. Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W.M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
@@ Appendix A Boundary Effects in the Lightcone
Because of the finite size of the cubes that are used to construct the lightcones (see Section 4.1 for details), there exist boundary effects at the interface between cubes that have the potential to cause under-richness in structures whose centers lie near the edge of the cubes. In addition, the lightcone is constructed in such a way that it subtends a circular region of radius of 2.3 deg, which causes additional boundary effects at the border of the lightcone along the transverse dimensions. In order to determine which structures are affected by these boundary effects, we utilize a version of the lightcone that probes 2 deeper than the version of lightcone that is used for the vast majority of our mocks (i.e., IRAC125.5 versus IRAC124.8) in order to search for simulated protostructures that are significantly under-rich in member galaxies relative to the population average. A deeper IRAC1 cut is used for this exercise in order to differentiate protostructures that are under-rich due to border effects from those that appear under-rich in the IRAC124.8 due to a shift in the luminosity function of member galaxies to fainter luminosities for astrophysical reasons.
As discussed in Section 4.3, we break the sample of simulated protostructures into redshift bins of over the redshift range and perform a (log) linear fit of versus with iterative clipping. All protostructures that are offset from the fit relation by 3 on the low end of this mass–richness relation are removed from our sample and are not considered in any of the metrics presented in this paper. More specifically, we do not use any subsumed structures that contain a clipped protostructure in the mass fit. Formally, we should also remove these structures from consideration in the estimate of the completeness and the purity calculation as well, but the inclusion of these structures have a negligible effect on the overall completeness numbers and would introduce a great deal of complexity into the purity calculation for likely the same result.
In total, the clipped protostructures comprise only 0.6% of the total protostructure catalog. Note that this process will also clip protostructures that are deficient in galaxies brighter than IRAC125.5, which will potentially induce some bias in our results, as, e.g., our completeness numbers may be overestimated for massive protostructures that are observed to be under-rich in galaxies (e.g., Newman et al., 2022). However, such protostructures appear, at least in our lightcones, to be rare (see also the simulation results presented in Newman et al. 2022). Figure 16 shows the –richness relation for all simulated protostructures as well as those that are identified as artificially under-rich. Also plotted are the R.A., decl., and redshift histogram of all simulated protostructures along with those that are identified as artificially under-rich.